# Relative periodic solutions in spatial Kepler problem with symmetric perturbation

### v1

@@ Abstract.
We study the spatial Kepler problem under a perturbation satisfying both rotational symmetry and reflection symmetry with respect to a plane perpendicular to the rotational axis. By applying recent results from [5] combined with Franks Theorem, we prove the existence of infinitely many relative periodic solutions contained in compact energy surface when the perturbation is sufficiently small and certain technical conditions hold. In addition, we also demonstrate the existence of a unique $z$-symmetric brake orbit which forms a hopf link with a plane relative periodic solutions contained in compact energy surface only requires the perturbation sufficiently small. Our results apply to the motion of a satellite around an ellipsoid with uniform mass distribution and to the $n$-pyramidal problem with one point mass moving along $z$-axis and other $n$ equal masses forming a regular $n$-gon perpendicular to the $z$-axis at all times.
This work is supported by the National Key R&D Program of China (2020YFA0713303). The first and second author is also supported by NSFC (No.12071255). The last author is also supported by NSFC (No. 12171253), Nankai Zhide Fundation and the Fundamental Research Funds for the Central Universities.
@@ 1. Introduction
where $(r,\theta,z)$ is the cylindrical coordinates of $\mathbb{R}^{3}$.
In celestial mechanics, one usually assumes that the celestial bodies are point masses without any size. Such an assumption is justified by the fact that when a celestial body is a perfect sphere with uniform mass distribution, outside the region of such a sphere, its gravitational force is equivalent to a point mass with all the masses concentrated at the center of the sphere.
Meanwhile a planet (or a star) is usually not a perfect sphere, but closer to an ellipsoid with small eccentricity. In this case, the gravitational force can be seen as a Kepler problem with a symmetric perturbation satisfying conditions (2). To be more precise, let’s assume the center of the ellipsoid is located at the origin, and under the spherical coordinates,
@@
Here $R$ denotes the mean radius, $\mathfrak{e}$ the eccentricity, and $P_{2}$ the second Legendre polynomial. For $\mathfrak{e}>0$, the spheroid is oblate (flattened along the $z$-axis), while for $\mathfrak{e}<0$, it is prolate (elongated along the $z$-axis). See [6] for detailed discussions.
@@
forms a compact manifold homeomorphic to $\mathbb{S}^{3}$. An orbit is called a relative periodic solution if it becomes periodic modulo the $\theta$-coordinate. Our main result states:
@@
a three-dimensional manifold. This configuration exhibits the following fundamental properties,
@@
For any $\varepsilon\geq 0$ small enough, there is a continuous family of subset $\Gamma_{\varepsilon}\subset\mathbb{R}\setminus\{0\}\times\mathbb{R}$ with
@@
    $\mathcal{M}(h,\varpi,\varepsilon)$ forms a compact contact manifold homeomorphic to $\mathbb{S}^{3}$ endowed with contact form $\lambda$.
@@
    There exists $R$ close to $\varpi^{2}/2$ satisfying $\min\Pi_{r}\mathcal{H}(h,\varpi,\varepsilon)>R$, where $\Pi_{r}$ denotes projection to the $r$-axis and
@@
Here $\mathcal{H}(h,\varpi,\varepsilon)$ represents the Hill’s region. In most of the previous studies of the perturbed Kepler problem, the collision singularity is usually regularized following Levi-Civita [14], Kustaanheimo-Stiefel [13] or Moser [19]. As a result, it is unclear whether the obtained solutions containing collision or not, for examples in [4] and [22]. Meanwhile in our case, the ellipsoid problem is not a point mass. The avoidance of collision is crucial. Property (b) guarantees orbital avoidance of the central body for sufficiently large angular momentum $\varpi$.
This proposition enables the application of symplectic and contact geometric methods to analyze mechanical systems. Recent developments in this direction include notable results for the restricted three-body problem [2, 1, 18] and isosceles three-body problem [9].
The foundational concept of global surfaces of section originated with Poincaré’s work [21] on the circular planar restricted three-body problem. His fundamental theorem on area-preserving twist maps of the annulus, later proved by Birkhoff [3], became known as the Poincaré-Birkhoff theorem. Franks [7] later generalized this result, showing that the existence of one periodic point implies infinitely many periodic points on the annulus. This framework extends to contact manifolds $(\mathcal{M},\lambda)$ homeomorphic to $\mathbb{S}^{3}$: If the Reeb flow admits a disk-like global surface of section, the system must contain either two or infinitely many periodic orbits. Recent breakthrough work by Cristofaro-Gardiner, Hryniewicz, Hutchings, and Liu [5] revealed an algebraic relationship between the contact volume and the action and Seifert rotation number of Reeb orbits (defined in Sections 2 and 3) for contact manifolds with exactly two Reeb orbits. Crucially, violation of this relationship guarantees the existence of infinitely many periodic orbits. In our setting, the reflection symmetry condition in (2) ensures a special periodic solution $\xi_{\varepsilon}(t)$ entirely contained in the $(p_{r},r)$-plane that bounds a global surface of section (see Section 3). Through calculation of this orbit’s action and Seifert rotation number combined with the algebraic relationship, we obtain:
@@
For sufficiently small $\varepsilon$ and parameters $(\varpi,h)\in\Gamma_{\varepsilon}$, the reduced Hamiltonian system (6) admits infinitely many periodic orbits when one of the conditions from Theorem 3.4 is satisfied.
While the above result proves the existence of infinitely many periodic orbits, it requires some additional conditions. Meanwhile we can show the existence of at least one special periodic orbit, as long as the perturbation is small enough. To explain it, we define two types of periodic solutions as below.
@@
  1. (a).
    is a brake orbit, if there are two different time moments $t_{1}\neq t_{2}$, such that the momentum variables all vanish, i.e., $p_{r}(t_{i})=p_{z}(t_{i})=0$, $i=1,2$;
  2. (b).
    is a $z$-symmetric orbit, if $p_{r}(t_{0})=z(t_{0})=0$ for some $t_{0}\in\mathbb{R}$ and
@@
A brake orbit must be a periodic solution and when it is also $z$-symmetric, we say it is a $z$-symmetric brake orbit.
@@
Our results extend beyond the ellipsoid problem to a special $(1+n)$-body configuration called the $n$-pyramidal problem. After fixing the center of mass at the origin, there is an invariant sub-system, where one of the masses $m_{0}$ always moving on the $z$-axis and the remaining $n$ equal masses $m$ always forming a regular $n$-gon with the plane containing them always perpendicular to the $z$-axis. Then the movement of one of the mass formed the $n$-gon can represent the movement of whole system. Let $\varepsilon=m/m_{0}$ denote the mass ratio, and fix energy $h<0$ with $z$-axis angular momentum $\varpi\neq 0$ such that the energy surface $\mathcal{M}(h,\varpi,\varepsilon)$ becomes compact, then we have the following Theorem.
@@
The $n$-pyramidal problem admits infinitely many relative periodic solutions within $\mathcal{M}(h,\varpi,\varepsilon)$ when the mass ratio $\varepsilon\in(0,\varepsilon_{0}(n))$ with sufficiently small $\varepsilon_{0}(n)$ and $2\leq n\leq 472$.
@@
When $n=2$, the $n$-pyramidal problem is the well-known isosceles three body problem, which has been studied by many authors. Moeckel [16] characterized collision asymptotics in this system. Using Maslov-type index (also called Conley-Zehnder index) and contact geometry, the first and third authors and their collaborators [9] established a criterion of existence of infinitely many relative periodic orbits. Grotta-Ragazzo, Liu and Salomão [8] obtained infinitely many relative periodic orbits when the eccentricity of Euler orbit is small enough using Birkhoff normal form. Our theorem solves the problem for eccentricity of Euler orbit $e\in(0,1)$ but mass ratio $\varepsilon$ is small enough.
@@
For techinical reason, the above theorem does not apply when $n>473$. However we believe the same result should still hold.
The paper is organized as follows. In Section 2, we introduce reduced Kepler problem and obtain some results which is needed in the perturbed cases. In Section 3, we consider the perturbed system. Theorem 1.2 and Theorem 1.3 are proved. In Section 4, we introduce the ellipsoid problem and prove Theorem 1.1. In Section 5, we introduce the $n$-pyramidal problem and prove Theorem 1.4.
@@ 2. The Kepler Problem
In this section, we obtain some results of the Kepler problem that will be needed in the perturbed cases. Recall that in Cartesian coordinates $q=(x,y,z)\in\mathbb{R}^{3}$ with the corresponding momentum variable $p=(\dot{x},\dot{y},\dot{z})\in\mathbb{R}^{3}$. The Hamiltonian of the Kepler problem can be written as
@@
Meanwhile in cylindrical coordinates $(r,\theta,z)$ with $(p_{r},p_{\theta},p_{z})=(\dot{r},r^{2}\dot{\theta},\dot{z})$, the above Hamiltonian can be written correspondingly as
@@
which is a system with two degrees of freedom. If $h$ is the energy constant, we get the 3 dimensional energy surface
@@
Let $\Pi_{r,z}$ denote its projection to the $(r,z)$-plane, we have the Hill’s region as below
@@
    if $1<2h\varpi^{2}<0$, then $\mathcal{M}(h,\varpi)$ is a compact manifold homeomorphic to $\mathbb{S}^{3}$ with
@@
$\nabla H_{\varpi}=\left(p_{r},p_{z},-\frac{\varpi^{2}}{r^{3}}+\frac{r}{(r^{2}+z^{2})^{3/2}},\frac{z}{(r^{2}+z^{2})^{3/2}}\right).$
@@
Therefor $\mathcal{M}(h,\varpi)$ is bounded in $z$ direction if and only if
@@
Then $\mathcal{M}(h,\varpi)$ is unbounded, when $2h\varpi^{2}>0$. This proves property (d).
For property (c), we only need to show $\mathcal{M}(h,\varpi)$ is bounded in $r$ direction, when $1<2h\varpi^{2}<0$. Consider the boundary of $\mathcal{H}(h,\varpi)$:
@@
Because of $V_{\varpi}(r,z)$ is increasing with respect to $z^{2}$, so we only need to see $z=0$. Solve the equation
@@
we have $r^{\pm}=\frac{\varpi^{2}}{1\pm\sqrt{1+2h\varpi^{2}}}>\frac{\varpi^{2}}{2}$. So we have $\Pi_{r}\mathcal{H}(h,\varpi)=[r^{+},r^{-}]$ is bounded. Therefor $\mathcal{M}(h,\varpi)$ is compact. By Morse theory, $\mathcal{M}(h,\varpi)$ is homeomorphic to $\mathbb{S}^{3}$.
For simplicity, we denote $\mathcal{M}(h,\varpi)$ and $\mathcal{H}(h,\varpi)$ by $\mathcal{M}$ and $\mathcal{H}$ respectively below, when there is no confusion. Without loss of generality, let $\varpi>0$ in the rest of this section.
The Hamiltonian vector field $X_{H_{\varpi}}$ is determined by $\omega(\cdot,X_{H_{\varpi}})=\mathrm{d}H_{\varpi}$, where $\omega=\mathrm{d}p_{r}\wedge\mathrm{d}r+\mathrm{d}p_{z}\wedge\mathrm{d}z$ is the symplectic form. Since $H_{\varpi}$ is a mechanical Hamiltonian, the hypersurface $\mathcal{M}$ is of contact type. That is there exists a 1-form $\lambda$ on $\mathcal{M}$ satisfying $\lambda\wedge\mathrm{d}\lambda\neq 0$, so that its Reeb vector field $R$, determined by $\mathrm{d}(R,\cdot)\equiv 0$ and $\lambda(R)=1$, is parallel to $X_{H_{\varpi}}$ and $\mathrm{d}\lambda=\omega$.
From now on we will focus on the cases with $2h\varpi^{2}\in(-1,0)$. By Proposition 2.1, $\mathcal{M}$ is a compact manifold homeomorphic to $\mathbb{S}^{3}$. Following the definitions in [5], we define its contact volume as
@@
and using Stokes formula
@@
where $S$ is the region bounded by $\mathcal{M}$, i.e. $\partial S=\mathcal{M}$.
@@
It corresponds to an elliptic orbit of the Kepler problem with eccentrity $e$ and entirely contained in the $(x,y)$-plane. As a result, its action value is
@@
where $\Upsilon$ is the region contained in the $(p_{r},r)$-plane and bounded by $\xi_{0}(t)$, i.e. $\partial\Upsilon=\xi_{0}(t)$.
@@
(21) $\int_{-\pi/2}^{\pi/2}\frac{\cos^{2}\theta}{\cos^{2}\theta+a\sin^{2}\theta}\mathrm{d}\theta$ $=\int_{-\pi/2}^{\pi/2}\frac{1}{1+a\tan^{2}\theta}\mathrm{d}\theta=\int_{-\infty}^{+\infty}\frac{1}{1+at^{2}}\cdot\frac{1}{1+t^{2}}\mathrm{d}t$
@@
(22) $\int_{-\pi/2}^{\pi/2}\frac{\cos^{2}\theta}{1+a\sin\theta}\mathrm{d}\theta$ $=\int_{-\pi/2}^{\pi/2}\frac{\cos^{2}\theta}{1-a^{2}\sin^{2}\theta}\mathrm{d}\theta-\int_{-\pi/2}^{\pi/2}\frac{a\cos^{2}\theta\sin\theta}{1-a^{2}\sin^{2}\theta}\mathrm{d}\theta$
$=\int_{-\pi/2}^{\pi/2}\frac{\cos^{2}\theta}{\cos^{2}\theta+(1-a^{2})\sin^{2}\theta}\mathrm{d}\theta=\frac{\pi}{1+\sqrt{1-a^{2}}}.$
Since $\cos^{2}\varphi^{+}=-2h\varpi^{2}=1-e^{2}$ and $\sin^{2}\varphi^{+}=1+2h\varpi^{2}=e^{2}$, we choose $\sin\varphi=e\sin\theta$, we conclude that $\cos\varphi\mathrm{d}\varphi=e\cos\theta\mathrm{d}\theta$. we compute that
@@
$=\frac{e^{2}}{-h}\int_{-\pi/2}^{\pi/2}\frac{\cos^{2}\theta}{\cos^{2}\theta+(1-e^{2})\sin^{2}\theta}\mathrm{d}\theta=\frac{e^{2}}{-h}\cdot\frac{\pi}{1+\sqrt{1-e^{2}}}$
@@
$=2\sqrt{1-e^{2}}\int_{u^{-}}^{u^{+}}\frac{\sqrt{e^{2}/(4h^{2})-u^{2}}}{u+1/(-2h)}\mathrm{d}u=\sqrt{1-e^{2}}\cdot\frac{e^{2}}{-h}\int_{-\pi/2}^{\pi/2}\frac{\cos^{2}\theta}{1+e\sin\theta}\mathrm{d}\theta$
@@
The above computation imply $\text{Vol}(\mathcal{M})=A^{2}(\xi_{0}(t))$. ∎
@@
Let $X$ be a non-vanishing vector field on $\mathbb{S}^{3}$ with flow $\varphi_{t}^{X}$. A disk-like global surface of section is an embedded closed 2-dim disk $D\subset\mathbb{S}^{3}$ satisfying
@@
When $2h\varpi^{2}\in(-1,0)$, $\Sigma_{\pm}$ given below, are two disk-like global surfaces of section of the Hamiltonian flow $\varphi_{t}$ corresponding to the Hamiltonian vector field $X_{H_{\varpi}}$,
@@
The boundary $\partial\Sigma_{\pm}=\{z=0,p_{z}=0\}\cap\mathcal{M}(h,\varpi)$ coincides with the orbit $\xi_{0}(t)$, which is tangent to $X_{H_{\varpi}}$.
@@
Consider a trajectory $\xi(t)=(p_{r}(t),p_{z}(t),r(t),z(t))$ starting from $\Sigma_{+}^{\circ}$. Suppose no $t_{1}>0$ exists with $p_{z}(t_{1})=0$ and $z(t_{1})>0$. Then $z(t)$ monotonically increases to some $z_{0}>0$, and the $\omega$-limit set of $\xi(t)$ lies in $\{z=z_{0}\}\cap\mathcal{M}$. By invariance of $\omega$-limit set under $\varphi_{t}$, there exists a solution $\hat{\xi}(t)=(\hat{p}_{r}(t),\hat{p}_{z}(t),\hat{r}(t),\hat{z}(t))$ with $\hat{z}(t)=z_{0}>0$ and $\hat{p}_{z}(t)=\dot{\hat{z}}(t)=0$. However,
$\dot{\hat{p}}_{z}(t)=\ddot{\hat{z}}(t)=-\frac{\hat{z}(t)}{(\hat{r}^{2}(t)+\hat{z}^{2}(t))^{3/2}}<0,$
@@
where $\Omega=q\times\dot{q}=(\varpi_{1},\varpi_{2},\varpi_{3})\in\mathbb{R}^{3}$ is the angular momentum vector, $\hat{e}=\sqrt{1+2h|\Omega|^{2}}$ is the eccentricity, $\hat{\theta}(t)$ satisfies $\rho^{2}(t)\mathrm{d}\hat{\theta}/\mathrm{d}t=|\Omega|$ and $\hat{\theta}(0)=0$ is the true anomaly and $g$ is the argument of perigee. The third component $\varpi=\varpi_{3}$ corresponds to the angular momentum about the $z$-axis.

### v2

@@ Abstract.
The spatial Kepler problem with a perturbation satisfying the rotational symmetry w.r.t. the $z$-axis and the reflection symmetry w.r.t. the $(x,y)$-plane, can be reduced to an Hamiltonian system with 2 degrees of freedom after fixing the angular momentum. For small enough perturbations, we show that for certain choices of energy and angular momentum, the corresponding energy surface is compact and diffeomorphic to $\mathbb{S}^{3}$, and on each compact energy surface there is a unique $z$-symmetric brake orbit, which forms a Hopf link with a planar relative periodic orbit. Moreover under some additional technical assumptions, by applying recent results from symplectic dynamics (Cristofaro-Gardiner et al 2023 Geom. Topol. 27 3801–31) and Franks’ Theorem, we prove there are infinitely many relative periodic orbits on each compact energy surface. These results can be applied to the motion of a satellite around a uniformly mass-distributed ellipsoid and the $n$-pyramidal problem, where one point mass moves along the $z$-axis and $n$ other equal point masses form a regular $n$-gon perpendicular to the $z$-axis.
This work is supported by the National Key R&D Program of China (2020YFA0713303). The first author is also supported by the National Natural Science Foundation of China (No. 12521001) and Taishan Scholars Climbing Program of Shandong (TSPD20240802). The last author is also supported by NSFC (No. 12171253), Nankai Zhide Fundation and the Fundamental Research Funds for the Central Universities.
* Corresponding author: Guowei Yu
@@ 1. Introduction
where $(r,\theta,z)$ are the cylindrical coordinates of $\mathbb{R}^{3}$.
In celestial mechanics, one usually assumes that the celestial bodies are point masses without any size. Such an assumption is justified by the fact that when a celestial body is a perfect sphere with uniform mass distribution, outside the region of such a sphere, its gravitational force is equivalent to a point mass with all the mass concentrated at the center of the sphere.
Meanwhile a planet (or a star) is usually not a perfect sphere, but closer to an ellipsoid with a small eccentricity. In this case, the gravitational force can be seen as a Kepler problem with a symmetric perturbation satisfying conditions (2). To be more precise, let’s assume the center of the ellipsoid is located at the origin, and in spherical coordinates,
@@
Here $R$ denotes the mean radius, $\mathfrak{e}$ the eccentricity, and $P_{2}$ the second Legendre polynomial. For $\mathfrak{e}>0$, the spheroid is oblate (flattened along the $z$-axis), while for $\mathfrak{e}<0$, it is prolate (elongated along the $z$-axis), for details see [6, page 29-33].
@@
forms a compact manifold diffeomorphic to $\mathbb{S}^{3}$. An orbit is called a relative periodic solution if it becomes periodic modulo the $\theta$-coordinate. Our main result states:
@@
a three-dimensional manifold. This configuration exhibits the following fundamental properties
@@
For any $\varepsilon\geq 0$ small enough, there is a continuous family of subsets $\Gamma_{\varepsilon}\subset\mathbb{R}\setminus\{0\}\times\mathbb{R}$ with
@@
    $\mathcal{M}(h,\varpi,\varepsilon)$ forms a compact contact manifold diffeomorphic to $\mathbb{S}^{3}$ endowed with contact form $\lambda$.
@@
    There exists $R$ close to $\varpi^{2}/2$ satisfying $\min\Pi_{r}\mathcal{H}(h,\varpi,\varepsilon)>R$, where $\Pi_{r}$ denotes the projection to the $r$-axis and
@@
Here $\mathcal{H}(h,\varpi,\varepsilon)$ represents Hill’s region. In most of the previous studies of the perturbed Kepler problem, the collision singularity is usually regularized following Levi-Civita [14], Kustaanheimo-Stiefel [13] or Moser [19]. As a result, it is unclear whether the obtained solutions contain collisions or not, for example in [4] and [22]. Meanwhile, in our case, the central body in the ellipsoid problem is not a point mass. The avoidance of collisions is crucial. Property (b) guarantees orbital avoidance of the central body for sufficiently large angular momentum $\varpi$.
This proposition enables the application of symplectic and contact geometric methods to analyze mechanical systems. Recent developments in this direction include notable results for the restricted three-body problem [2, 1, 18] and the isosceles three-body problem [9].
The foundational concept of global surfaces of section originated with Poincaré’s work [21] on the circular planar restricted three-body problem. His fundamental theorem on area-preserving twist maps of the annulus, later proved by Birkhoff [3], became known as the Poincaré-Birkhoff Theorem. Franks [7] later generalized this result, showing that the existence of one periodic point implies infinitely many periodic points on the annulus. This framework extends to contact manifolds $(\mathcal{M},\lambda)$ diffeomorphic to $\mathbb{S}^{3}$: If the Reeb flow admits a disk-like global surface of section, the system must contain either two or infinitely many periodic orbits. Recent breakthrough work by Cristofaro-Gardiner, Hryniewicz, Hutchings, and Liu [5] revealed an algebraic relationship between the contact volume and the action and Seifert rotation number of Reeb orbits (defined in Sections 2 and 3) for contact manifolds with exactly two Reeb orbits. Crucially, violation of this relationship guarantees the existence of infinitely many periodic orbits. In our setting, the reflection symmetry condition in (2) ensures a special periodic solution $\xi_{\varepsilon}(t)$ entirely contained in the $(p_{r},r)$-plane that bounds a global surface of section (see Section 3). Through the calculations of this orbit’s action and Seifert rotation number combined with the algebraic relationship, we obtain:
@@
For sufficiently small $\varepsilon$ and parameters $(\varpi,h)\in\Gamma_{\varepsilon}$, the reduced Hamiltonian system (6) admits infinitely many periodic orbits when one of the two conditions in Theorem 3.4 is satisfied.
While the above result proves the existence of infinitely many periodic orbits, it requires some additional conditions. Meanwhile, we can show the existence of at least one special periodic orbit as long as the perturbation is small enough. To explain this, we define two types of periodic solutions as follows.
@@
  1. (a)
    is a brake orbit if there exist two distinct time moments $t_{1}\neq t_{2}$, such that the momentum variables all vanish, i.e., $p_{r}(t_{i})=p_{z}(t_{i})=0$, for $i=1,2$;
  2. (b)
    is a $z$-symmetric orbit if $p_{r}(t_{0})=z(t_{0})=0$ for some $t_{0}\in\mathbb{R}$ and
@@
A brake orbit must be a periodic solution, and when it is also $z$-symmetric, we say it is a $z$-symmetric brake orbit.
@@
Our results extend beyond the ellipsoid problem to a special $(1+n)$-body configuration called the $n$-pyramidal problem. After fixing the center of mass at the origin, there is an invariant subsystem, where one of the masses $m_{0}$ always moves on the $z$-axis and the remaining $n$ equal masses $m$ always form a regular $n$-gon with the plane containing them always perpendicular to the $z$-axis. Then the movement of one of the masses forming the $n$-gon can represent the movement of the whole system. Let $\varepsilon=m/m_{0}$ denote the mass ratio. Then for energy $h<0$ and $z$-axis angular momentum $\varpi\neq 0$, such that the corresponding energy surface $\mathcal{M}(h,\varpi,\varepsilon)$ is compact, we have the following theorem.
@@
The $n$-pyramidal problem admits infinitely many relative periodic solutions within $\mathcal{M}(h,\varpi,\varepsilon)$, when the mass ratio $\varepsilon\in(0,\varepsilon_{0}(n))$ with sufficiently small $\varepsilon_{0}(n)$ and $2\leq n\leq 472$.
@@
When $n=2$, the $n$-pyramidal problem is the well-known isosceles three body problem, which has been studied by many authors. Moeckel [16] characterized collision asymptotics in this system. Using Maslov-type index (also called Conley-Zehnder index) and contact geometry, the first and third authors and their collaborators [9] established a criterion for the existence of infinitely many relative periodic orbits. Grotta-Ragazzo, Liu and Salomão [8] obtained infinitely many relative periodic orbits when the eccentricity of Euler orbit is small enough using Birkhoff normal form. Our theorem solves the problem for the eccentricity of the Euler orbit $e\in(0,1)$ but the mass ratio $\varepsilon$ is small enough.
@@
For technical reasons, the above theorem does not apply when $n>472$. However, we believe the same result should still hold.
The paper is organized as follows. In Section 2, we introduce the reduced Kepler problem and obtain some results which are needed in the perturbed cases. In Section 3, the perturbed system is studied and proofs of Theorem 1.2 and Theorem 1.3 are given. In Section 4, we introduce the ellipsoid problem and prove Theorem 1.1. In Section 5, we introduce the $n$-pyramidal problem and prove Theorem 1.4.
@@ 2. The Kepler Problem
In this section, we obtain some results of the Kepler problem that will be needed in the perturbed cases. Recall that in Cartesian coordinate $q=(x,y,z)\in\mathbb{R}^{3}$ with the corresponding momentum variable $p=(\dot{x},\dot{y},\dot{z})\in\mathbb{R}^{3}$, the Hamiltonian of the Kepler problem can be written as
@@
Meanwhile, in cylindrical coordinates $(r,\theta,z)$ with $(p_{r},p_{\theta},p_{z})=(\dot{r},r^{2}\dot{\theta},\dot{z})$, the above Hamiltonian can be written correspondingly as
@@
which is a system with two degrees of freedom. If $h$ is the energy constant, we get the three-dimensional energy surface
@@
Let $\Pi_{r,z}$ denote the projection to the $(r,z)$-plane, we have Hill’s region as follows:
@@
    if $1<2h\varpi^{2}<0$, then $\mathcal{M}(h,\varpi)$ is a compact manifold diffeomorphic to $\mathbb{S}^{3}$ with
@@
[…]
@@
Therefore $\mathcal{M}(h,\varpi)$ is bounded in $z$-direction if and only if
@@
Then $\mathcal{M}(h,\varpi)$ is unbounded when $2h\varpi^{2}>0$. This proves property (d).
For property (c), we only need to show $\mathcal{M}(h,\varpi)$ is bounded in $r$-direction, when $1<2h\varpi^{2}<0$. Consider the boundary of $\mathcal{H}(h,\varpi)$:
@@
Because $V_{\varpi}(r,z)$ is increasing with respect to $z^{2}$, so we only need to consider $z=0$. Solving the equation
@@
we have $r^{\pm}=\frac{\varpi^{2}}{1\pm\sqrt{1+2h\varpi^{2}}}>\frac{\varpi^{2}}{2}$. So we have $\Pi_{r}\mathcal{H}(h,\varpi)=[r^{+},r^{-}]$, which is bounded. Therefore $\mathcal{M}(h,\varpi)$ is compact. Because the unique critical point $(0,0,\varpi^{2},0)$ of $H_{\varpi}$ is nondegenerate, when $h$ is slightly greater than $\frac{1}{2\varpi^{2}}$, the energy surface $\mathcal{M}(h,\varpi)$ is diffeomorphic to $\mathbb{S}^{3}$. When $1<2h\varpi^{2}<0$, the regular energy surfaces $\mathcal{M}(h,\varpi)$ are diffeomorphic to one another. We can also use the Reeb sphere theorem to prove $\mathcal{M}(h,\varpi)$ is diffeomorphic to $\mathbb{S}^{3}$. The Morse function $f(p_{r},p_{z},r,z)=r$ has exactly two critical points $(0,0,r^{+},0)$ and $(0,0,r^{-},0)$ on $\mathcal{M}(h,\varpi)$. Combining together with the uniqueness of the smooth structure on $\mathcal{M}(h,\varpi)$, $\mathcal{M}(h,\varpi)$ is diffeomorphic to $\mathbb{S}^{3}$.
For simplicity, we denote $\mathcal{M}(h,\varpi)$ and $\mathcal{H}(h,\varpi)$ by $\mathcal{M}$ and $\mathcal{H}$ respectively below, when there is no confusion. Without loss of generality, let’s assume $\varpi>0$ in the rest of this section.
The Hamiltonian vector field $X_{H_{\varpi}}$ is determined by $\omega(\cdot,X_{H_{\varpi}})=\mathrm{d}H_{\varpi}$, where $\omega=\mathrm{d}p_{r}\wedge\mathrm{d}r+\mathrm{d}p_{z}\wedge\mathrm{d}z$ is the symplectic form. Since $H_{\varpi}$ is a mechanical Hamiltonian, the hypersurface $\mathcal{M}$ is of contact type. That is, there exists a 1-form $\lambda$ on $\mathcal{M}$ satisfying $\lambda\wedge\mathrm{d}\lambda\neq 0$, so that its Reeb vector field $R$, determined by $\mathrm{d}\lambda(R,\cdot)\equiv 0$ and $\lambda(R)=1$, is parallel to $X_{H_{\varpi}}$ and $\mathrm{d}\lambda=\omega$.
From now on we will focus on the cases with $2h\varpi^{2}\in(-1,0)$. By Proposition 2.1, $\mathcal{M}$ is a compact manifold diffeomorphic to $\mathbb{S}^{3}$. Following the definitions in [5], we define its contact volume as
@@
and using Stokes’ formula
@@
where $S$ is the region bounded by $\mathcal{M}$, i.e., $\partial S=\mathcal{M}$.
@@
It corresponds to an elliptic orbit of the Kepler problem with eccentricity $e$ and entirely contained in the $(x,y)$-plane. As a result, its action value is
@@
where $\Upsilon$ is the region contained in the $(p_{r},r)$-plane and bounded by $\xi_{0}(t)$, i.e., $\partial\Upsilon=\xi_{0}(t)$.
@@
[…]
@@
[…]
$2.\displaystyle=\int_{-\pi/\penalty 502}^{\pi/\penalty 502}\frac{\cos^{2}\theta}{\cos^{2}\theta+(1-a^{2})\sin^{2}\theta}\mathrm{d}\theta=\frac{\pi}{1+\sqrt{1-a^{2}}}$.
Since $\cos^{2}\varphi^{+}=-2h\varpi^{2}=1-e^{2}$ and $\sin^{2}\varphi^{+}=1+2h\varpi^{2}=e^{2}$, we choose $\sin\varphi=e\sin\theta$, and we conclude that $\cos\varphi\mathrm{d}\varphi=e\cos\theta\mathrm{d}\theta$. We compute that
@@
$2\displaystyle=\frac{e^{2}}{-h}\int_{-\pi/\penalty 502}^{\pi/\penalty 502}\frac{\cos^{2}\theta}{\cos^{2}\theta+(1-e^{2})\sin^{2}\theta}\mathrm{d}\theta=\frac{e^{2}}{-h}\cdot\frac{\pi}{1+\sqrt{1-e^{2}}}$
@@
[…]
@@
The above computation implies $\text{Vol}(\mathcal{M})=A^{2}(\xi_{0}(t))$. ∎
@@
Let $X$ be a non-vanishing vector field on $\mathbb{S}^{3}$ with flow $\varphi_{t}^{X}$. A disk-like global surface of section is an embedded closed 2-dimensional disk $D\subset\mathbb{S}^{3}$ satisfying
@@
When $2h\varpi^{2}\in(-1,0)$, $\Sigma_{\pm}$ given below are two disk-like global surfaces of section of the Hamiltonian flow $\varphi_{t}$ corresponding to the Hamiltonian vector field $X_{H_{\varpi}}$,
@@
The boundary $\partial\Sigma_{\pm}=\{z=0,p_{z}=0\}\cap\mathcal{M}(h,\varpi)$ coincide with the orbit $\xi_{0}(t)$, which is tangent to $X_{H_{\varpi}}$.
@@
Consider a trajectory $\xi(t)=(p_{r}(t),p_{z}(t),r(t),z(t))$ starting from $\Sigma_{+}^{\circ}$. Suppose no $t_{1}>0$ exists with $p_{z}(t_{1})=0$ and $z(t_{1})>0$. Then $z(t)$ monotonically increases to some $z_{0}>0$, and the $\omega$-limit set of $\xi(t)$ lies in $\{z=z_{0}\}\cap\mathcal{M}$. By the invariance of the $\omega$-limit set under $\varphi_{t}$, there exists a solution $\hat{\xi}(t)=(\hat{p}_{r}(t),\hat{p}_{z}(t),\hat{r}(t),\hat{z}(t))$ with $\hat{z}(t)=z_{0}>0$ and $\hat{p}_{z}(t)=\dot{\hat{z}}(t)=0$. However,
[…]
@@
where $\Omega=q\times\dot{q}=(\varpi_{1},\varpi_{2},\varpi_{3})\in\mathbb{R}^{3}$ is the angular momentum vector, $\hat{e}=\sqrt{1+2h|\Omega|^{2}}$ is the eccentricity, $\hat{\theta}(t)$ which satisfies $\rho^{2}(t)\mathrm{d}\hat{\theta}/\mathrm{d}t=|\Omega|$ and $\hat{\theta}(0)=0$ is the true anomaly and $g$ is the argument of perigee. The third component $\varpi=\varpi_{3}$ corresponds to the angular momentum about the $z$-axis.